# Discrete time crystal for periodic-field sensing with quantum-enhanced precision

### v1

@@ Abstract
Sensing periodic-fields using quantum sensors has been an active field of research. In many of these scenarios, the quantum state of the probe is flipped regularly by the application of $\pi$-pulses to accumulate information about the target periodic-field. The emergence of a discrete time crystalline phase, as a nonequilibrium phase of matter, naturally provides oscillations in a many-body system with an inherent controllable frequency. They benefit from long coherence time and robustness against imperfections, which makes them excellent potential quantum sensors. In this paper, through theoretical and numerical analysis, we show that a disorder-free discrete time crystal probe can reach the ultimate achievable precision for sensing a periodic-field. As the amplitude of the periodic-field increases, the discrete time crystalline order diminishes, and the performance of the probe decreases remarkably. Nevertheless, the obtained quantum enhancement in the discrete time crystal phase shows robustness against different imperfections in the protocol. Finally, we propose the implementation of our protocol in ultra-cold atoms in optical lattices.
@@ I Introduction
Metrological tasks can harness quantum features, such as superposition and entanglement, to achieve greater precision compared to their classical counterparts [1]. To estimate an unknown parameter $h_{a}$, the precision of the estimation $\delta h_{a}$, quantified by standard deviation, is fundamentally lower bounded by Cramér-Rao inequality [2, 3] as $\delta h_{a}{\geq}1/\sqrt{M\mathcal{F}_{Q}}$ where $M$ is the number of observations and $\mathcal{F}_{Q}$ is the quantum Fisher information (QFI) [4]. If the Hamiltonian $H{=}H_{0}{+}h_{a}H_{1}$, in which $H_{0}$ independently describes the physics of the probe and $H_{1}$ is related to the parameter, encodes $h_{a}$ through the unitary evolution $U(t){=}e^{-itH}$, it has been shown that the QFI is upper bounded by $\mathcal{F}_{Q}{\leq}t^{2}||H_{1}||^{2}$, where $||H_{1}||$ is the spectral norm of $H_{1}$ and can be related to the size of the probe $L$ as $||H_{1}||^{2}{\propto}L^{\beta}$ [5, 6, 7]. Although in the absence of quantum features, one obtains $\beta{=}1$, known as the standard quantum limit, their presence allows quantum-enhanced sensitivity, namely $\beta{>}1$. For $H_{0}{=}0$. This quantum enhancement can be achieved in phase shift detection by initializing the probe in an entangled Greenberger–Horne–Zeilinger (GHZ) state [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. However, this approach is limited by the difficulty in creating and maintaining such GHZ-type states, especially on a large scale [18, 19]. In or out of equilibrium, Strongly correlated many-body systems provide a valid alternative to this approach [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Due to $H_{0}{\neq}0$, many-body sensors benefit from the entanglement that is either intrinsically present in the spectrum or generated freely during the time evolution  [35].
Quantum sensors have also been developed for periodic field sensing inferring the amplitude [39, 40, 41], the frequency [42, 43] and the phase [44, 45, 46, 47, 48, 49] of oscillatory fields in the environment. In many such schemes, high-frequency spin echo pulses, typically used in dynamic decoupling [50, 51, 51, 52, 43, 53, 1] to overcome decoherence, are used to improve the accumulation of information about the time-varying signal in the quantum state of the probe. However, these high-frequency pulses often diminish spin correlations in the quantum state of the probe, which may decrease its sensing capability. Recent studies demonstrate that the spin correlations created by the system’s dynamics can be utilized to achieve enhanced periodic-field sensing [54, 55]. Among these proposals, systems with long-range spatiotemporal ordering, known as time crystals, are attracting increasing interest [56, 57, 58, 59, 60, 61].
@@
A well-known characteristic of DTCs is the formation of paired GHZ-type states with a $\pi$ quasienergy gap in their spectrum. The emergence of these large-scale entangled states that are responsible for subharmonic responses makes DTCs a promising candidate for quantum technological tasks. Thus far, DTCs have found applications in simulating complex systems [75], topologically protected quantum computation [114], quantum engine design [115], and metrology [59, 57, 61, 110]. Sensing in fully connected graphs [59] and gradient interacting systems [61] demonstrates the potential of DTCs for quantum-enhanced precision by achieving $\beta{=}3$ in sensing the exchange coupling While this surpasses the standard quantum limit, the ultimate limit determined by $||H_{1}||^{2}{\propto}L^{\beta}$ with $\beta{=}4$ in sensing the strength of two-body interactions, has not been obtained. Relying on the long coherence time of DTCs, theoretical [57] and experimental [110] results demonstrate the capability of DTCs for constructing highly frequency-selective sensors of AC magnetic fields. Nevertheless, effectively exploiting the inherent GHZ-type eigenstates in the spectrum of the DTCs is a challenge that has yet to be addressed.
In this work, through concrete theoretical and numerical analysis, we show that a devised probe in the DTC order can achieve ultimate estimation precision in a gradient periodic magnetic field sensing. The external periodic-field breaks site invariance symmetry and eliminates the DTC order with a rate that depends on the probe size. However, for small values of the periodic-field, our results demonstrate that in the DTC order, the probe can benefit from the inherent GHZ state in the spectrum and reach ultimate precision. The robustness of this performance has been analyzed against various types of defects and imperfections. We further provide details of our probe’s experimental implementation on ultra-cold atoms in optical lattices.
@@ II Quantum parameter estimation
In this section, we review the theory of quantum parameter estimation. This field seeks to determine an unknown parameter, $h_{a}$, encoded in the quantum state $\rho(h_{a})$. The key objective is to minimize the uncertainty in estimating $h_{a}$, represented by the standard deviation $\delta h_{a}$. This uncertainty is fundamentally limited by the Cramér-Rao inequality [2, 3], which relates the minimum error attainable by any locally unbiased estimator to the classical Fisher information (CFI) $\mathcal{F}_{C}(h_{a})$, as
@@
where $M$ represents the number of repetitions of the sensing protocol. For a given set of Positive Operator-Valued Measurement (POVM) operators $\{\Pi_{r}\}$, the probability of a measurement outcome $r$ is given by $p_{r}(h_{a}){=}{\rm Tr}[\rho(h_{a})\Pi_{r}]$. One can obtain the CFI as $\mathcal{F}_{C}(h_{a}){=}\sum_{r}p_{r}(h_{a})[\partial_{h_{a}}\ln p_{r}(h_{a})]^{2}$, with $\partial_{h_{a}}$ being the derivatives with respect to the parameter $h_{a}$. Optimizing over all possible POVMs $\{\Pi_{r}\}$ leads to the quantum Cramér-Rao inequality, which is bounded by the quantum Fisher information (QFI) $\mathcal{F}_{Q}(h_{a})$ as
@@
Thus, the QFI, as a central quantity in metrology, defines the maximum precision limit in quantum sensing. For pure quantum states $\rho(h_{a}){=}|\psi(h_{a})\rangle\langle\psi(h_{a})|$, the QFI takes a simple form as [4]
@@
The performance of a sensor is characterized by how the QFI scales with available resources, such as probe size and time. In particular, in closed systems, this relationship can be expressed as $\mathcal{F}_{Q}{\sim}n^{\alpha}L^{\beta}$, where $n$ and $L$ represent the duration time of the protocol and the size of the probe, respectively, with $\alpha$ and $\beta$ as the corresponding exponents. Without quantum effects, the QFI can scale at most quadratically with $n$ and linearly with $L$, represented as $\alpha{=}1$ and $\beta{=}1$, which is known as the standard limit. However, by exploiting quantum probes with quantum features, it is possible to enhance the scaling to $\alpha{>}1$ and $\beta{>}1$, a phenomenon called quantum-enhanced sensitivity. A notable case is when $\alpha{=}2$ and $\beta{=}2$, which is called Heisenberg scaling. The most common way to capture this scaling is by initializing a dynamical protocol in maximally entangled states, such as Greenberger–Horne–Zeilinger (GHZ) states, and encoding the desired parameter through a unitary evolution, while the interaction between the system’s components remains negligible. The limitations of this method prompt efforts to unlock the potential of many-body interacting systems. Strongly correlated many-body systems show a remarkable potential to go beyond the standard limit, with $\beta{>}1$, by harnessing a variety of quantum features such as criticality. Nonetheless, the obtainable precision is fundamentally bounded by the interaction details in the system [5, 6, 7].
@@ III Model
Figure 1: (a) Schematic for the DTC sensor introduced here. (b) Imbalance in the total magnetization of the chains $a$ and $b$ denoted as $\mathcal{I}(nT)$, versus the amplitude of the periodic-field $h_{a}$ at different stroboscopic times in a system of size $L{=}6$ with $\varepsilon{=}0.1$.
@@
shows persistent coherent oscillation between $\mathcal{I}(nT){=}1$ at even $n$’s obtained for $|\psi(0)\rangle$, and $\mathcal{I}(nT){=}{-}1$ at odd $n$’s obtained for the inversion partner of the initial state $|\tilde{\psi}(0)\rangle{=}\Pi_{j=1}^{L}\sigma^{ax}_{j}{\otimes}\sigma^{bx}_{j}|\psi(0)\rangle$. The origin of this period-doubling oscillations that last for $\tau{\propto}e^{\alpha L}$ with $\alpha{=}\ln(\varepsilon^{-1.033})$ is the emergence of $\pi$-Paired Floquet states as $|\Phi_{\pm}\rangle{\cong}(|\psi(0)\rangle{\pm}|\tilde{\psi}(0)\rangle){/}\sqrt{2}$ with eigenvalues differing by $\pi$ [111].
To investigate the sensing capability of this system, we expose the chain $a$ to a gradient periodic-field in the $z$-direction as
$H_{ac}(t){=}B_{a}(t)G^{az}{=}h_{a}\sin(2\pi f_{ac}t+\theta_{ac})\sum_{j=1}^{L}j\sigma_{j}^{az},$ (6)
where $h_{a}$, $f_{ac}$, and $\theta_{ac}{\in}[0,\pi/2)$ are amplitude, frequency, and the phase of the periodic-field, respectively. See Fig. 1 (a) for a system schematic. The cross-talk effect will be discussed later. The time propagator for one period cycle can be obtained as
[…]
@@
where $\mathcal{T}$ denotes the time-ordering operator. For the periodic-field that resonates with the binary quench, namely $T_{ac}{=}T$, one has $\Theta_{n}{=}\frac{(-1)^{n}h_{a}}{\pi J_{z}}\cos(\theta_{ac})$ for $t_{1}{=}t_{2}{=}1{/}2J_{z}$. Without loss generality, we set $J_{z}{=}1$, and $\theta_{ac}{=}0$. To show the effect of the implemented periodic-field on the DTC order, Fig. 1 (b) plots the imbalance $\mathcal{I}(nT)$ versus the amplitude of the periodic-field $h_{a}$ over tens of stroboscopic times $n$ when the system initialized in $|\psi(0)\rangle$. By increasing $h_{a}$,the system goes through a sharp phase transition from a stable DTC order to a region with no spontaneous breaking of DTTS in Eq. (4). Surprisingly, this transition, happening in a specific value denoted by $h_{a}{=}h_{a}^{\max}$ (dashed line), repeats in all the stroboscopic times, and, as we show latter, is robust against the deviation parameter $\varepsilon$.
@@
In the following, we first show that this phase transition can serve as a valuable resource for quantum-enhanced sensitivity and provide ultraprecise periodic-field sensing. Then, we complete our study by providing a robustness analysis of this enhancement to the perturbations in initialization, frequency offset, and crosstalk. We also explore the effect of $\varepsilon$ on the capability of our probe and propose an experimental protocol for its implementation.
@@ IV DTC sensor
Figure 2: (a) The QFI as a function of $h_{a}$ over different stroboscopic times. (b) The dynamical behavior of the QFI versus stroboscopic time $n$ at different values of $h_{a}$. In both panels (a) and (b), we set $L{=}7$ and $\varepsilon{=}0.1$. (c) QFI as a function of $h_{a}$ for systems of different sizes over $n{=}10$ period cycles when $\varepsilon{=}0.1$. (d) The values of QFI inside the DTC phase, namely $h_{a}{=}10^{-5}$, and in the transition point, namely $h_{a}{=}h_{a}^{\max}$, as a function of the system size $L$. The markers are the numerical results, and the solid lines provide the best fitting function as $\mathcal{F}_{Q}{\propto}L^{\beta}$ with $\beta{=}4.0902$ and $\beta{=}2.5887$ for a sensor in the DTC phase and in its transition point. The inset is the transition point $h_{a}^{\max}$ as a function of $L$. The best fitting function is obtained as $h_{a}^{\max}{\propto}L^{-1.0916}$.
[…]
$\partial_{h_{a}}U_{t}$ $={-i}U_{t}\sum_{k=1}^{n}\Big{(}\Pi_{l=1}^{k}U_{l}\Big{)}^{\dagger}\Big{(}\partial_{h_{a}}\Theta_{k}G^{az}\Big{)}\Big{(}\Pi_{l=1}^{k}U_{l}\Big{)}\leqslant iU_{t}G^{az}/\pi.$ (8)
@@
=$=\begin{cases}i(-1)^{k+1}G^{az}/\pi&(k-1)T<t_{k}\leqslant(k-1/2)T\\ i(-1)^{k+1}D_{k}^{az}/\pi&(k-1/2)T<t_{k}\leqslant kT\\ \end{cases}$ (10)
wherein $D_{k}^{az}{=}\sum_{j=1}^{L}j\alpha_{j}^{k}\sigma^{az}_{j}{\leqslant}\sum_{j=1}^{L}j\sigma^{az}_{j}{=}G^{az}$ and $\alpha_{j}{=}\Big{(}j^{2}\Theta_{k}^{2}+J_{ab}^{2}\cos^{2}(\sqrt{j^{2}\Theta_{k}^{2}+J_{ab}^{2}})\Big{)}/\Big{(}j^{2}\Theta_{k}^{2}+J_{ab}^{2}\Big{)}$. Using
@@
44 44 $\left(\langle\psi(0)|(\partial_{h_{a}}U_{t})^{\dagger}(\partial_{h_{a}}U_{t})|\psi(0)\rangle{-}|\langle\psi(0)|U_{t}^{\dagger}(\partial_{h_{a}}U_{t})|\psi(0)\rangle|^{2}\right)$
@@
and Eq. 8, one can simply obtain $\mathcal{F}_{Q}(t){\leqslant}4n^{2}\Delta G^{az}$ in which $\Delta G^{az}{=}\sqrt{\langle(G^{az})^{2}\rangle{-}\langle G^{az}\rangle^{2}}$ stands for the variance that is upper bounded by the square of the spectral norm defined as $||G^{az}||^{2}{=}(\lambda_{\max}{-}\lambda_{\min})^{2}$ with $\lambda_{\max}{=}L(L{-}1)/2$ and $\lambda_{\min}{=}{-}L(L{-}1)/2$ as the largest and smallest eigenvalues of $G^{az}$. This results in $\mathcal{F}_{Q}{\leqslant}4n^{2}L^{2}(L-1)^{2}.$ Capturing this ultimate precision in estimating $h_{a}$ requires satisfying two conditions. First, $\alpha_{j}{\simeq}1$, that requires $\varepsilon{=}0$ and $h_{a}{\ll}\sqrt{2}\pi^{2}/L$. Second, the probe is to be initialized in the maximally entangled state as $\frac{1}{\sqrt{2}}\big{(}|{\uparrow,\cdots,\uparrow}\rangle_{a}{\otimes}|{\downarrow,\cdots,\downarrow}\rangle_{b}{\pm}|{\downarrow,\cdots,\downarrow}\rangle_{a}{\otimes}|{\uparrow,\cdots,\uparrow}\rangle_{b}\big{)}$. In the following, we show that, although initialized in the product state $|\psi(0)\rangle$, achieving this ultimate precision even for $\varepsilon{\neq}0$ is permitted in our DTC probe.

### v2

@@ Abstract
Sensing periodic-fields using quantum sensors has been an active field of research. In many of these scenarios, the quantum state of the probe is flipped regularly by the application of $\pi$-pulses to accumulate information about the target periodic-field. The emergence of a discrete time crystalline phase, as a nonequilibrium phase of matter, naturally provides oscillations in a many-body system with an inherent controllable frequency. They benefit from long coherence time and robustness against imperfections, which makes them excellent potential quantum sensors. In this paper, through theoretical and numerical analysis, we show that a disorder-free discrete time crystal probe can reach the ultimate achievable precision for sensing a periodic-field. As the amplitude of the periodic-field increases, the discrete time crystalline order diminishes, and the performance of the probe decreases remarkably. Nevertheless, the obtained quantum enhancement in the discrete time crystal phase, which is experimentally accessible using standard projective measurements, shows robustness against different imperfections and dephasing noise in the protocol. Finally, we propose the implementation of our protocol in ultra-cold atoms in optical lattices.
@@ I Introduction
[…]
Quantum sensors have also been developed for periodic field sensing inferring the amplitude [39, 40, 41, 42], the frequency [43, 44] and the phase [45, 46, 47, 48, 49, 50] of oscillatory fields in the environment. In many such schemes, high-frequency spin echo pulses, typically used in dynamic decoupling [51, 52, 52, 53, 44, 54, 1] to overcome decoherence, are used to improve the accumulation of information about the time-varying signal in the quantum state of the probe. However, these high-frequency pulses often diminish spin correlations in the quantum state of the probe, which may decrease its sensing capability. Recent studies demonstrate that the spin correlations created by the system’s dynamics can be utilized to achieve enhanced periodic-field sensing [55, 56]. Among these proposals, systems with long-range spatiotemporal ordering, known as time crystals, are attracting increasing interest [57, 58, 59, 60, 61, 62].
@@
A well-known characteristic of DTCs is the formation of paired GHZ-type states with a $\pi$ quasienergy gap in their spectrum. The emergence of these large-scale entangled states that are responsible for subharmonic responses makes DTCs a promising candidate for quantum technological tasks. Thus far, DTCs have found applications in simulating complex systems [76], topologically protected quantum computation [115], quantum engine design [116], and metrology [60, 58, 62, 111, 117]. Sensing in fully connected graphs [60] and gradient interacting systems [62] demonstrates the potential of DTCs for quantum-enhanced precision by achieving $\beta{=}3$ in sensing the exchange coupling. While this surpasses the standard quantum limit, the ultimate limit determined by $H_{1}||^{2}{\propto}L^{\beta}$ with $\beta{=}4$ in sensing the strength of two-body interactions, has not been obtained. Relying on the long coherence time of DTCs, theoretical [58] and experimental [111] results demonstrate the capability of DTCs for constructing highly frequency-selective sensors of AC magnetic fields. Nevertheless, effectively exploiting the inherent GHZ-type eigenstates in the spectrum of the DTCs is a challenge that has yet to be addressed.
In this work, through concrete theoretical and numerical analysis, we show that a devised probe in the DTC order can achieve ultimate estimation precision in a gradient periodic magnetic field sensing. The external periodic-field breaks site invariance symmetry and eliminates the DTC order at a rate that depends on the probe size. However, for small values of the periodic-field, our results demonstrate that in the DTC order, the probe can benefit from the inherent GHZ state in the spectrum and achieve ultimate precision.
The robustness of this performance has been analyzed against various types of imperfections and dephasing noise. Interestingly, we find that in the sensing scheme, a simple configuration measurement can follow the scaling behavior of the QFI, concerning both time and size. We further provide details of our probe’s experimental implementation on ultra-cold atoms in optical lattices.
@@ II Quantum parameter estimation
In this section, we review the theory of quantum parameter estimation. The goal is to estimate an an unknown parameter, $h_{a}$, encoded in the quantum state $\rho(h_{a})$. The key objective is to minimize the uncertainty in estimating $h_{a}$, represented by the standard deviation $\delta h_{a}$. This uncertainty is fundamentally bounded by the Cramér-Rao inequality [2, 3], which relates the minimum error attainable by any locally unbiased estimator to the classical Fisher information (CFI) $\mathcal{F}_{C}(h_{a})$, as
@@
where $M$ represents the number of repetitions of the sensing protocol. For a given set of Positive Operator-Valued Measurement (POVM) operators $\{\Pi_{r}\}$, the probability of a measurement outcome $r$ is given by $p_{r}(h_{a}){=}{\rm Tr}[\rho(h_{a})\Pi_{r}]$. One can obtain the CFI as $\mathcal{F}_{C}(h_{a}){=}\sum_{r}p_{r}(h_{a})[\partial_{h_{a}}\ln p_{r}(h_{a})]^{2}$, with $\partial_{h_{a}}$ being the derivatives with respect to the parameter $h_{a}$. Optimizing over all possible POVMs $\{\Pi_{r}\}$ leads to the quantum Cramér-Rao inequality, which is bounded by the QFI $\mathcal{F}_{Q}(h_{a})$ as
@@
Thus, the QFI, as a central quantity in metrology, determines the ultimate precision limit in quantum sensing. For pure quantum states $\rho(h_{a}){=}|\psi(h_{a})\rangle\langle\psi(h_{a})|$, the QFI takes a simple form as [4]
@@
The performance of a sensor is characterized by how the QFI scales with available resources, such as probe size and time. In particular, in closed systems, this relationship can be expressed as $\mathcal{F}_{Q}{\sim}n^{\alpha}L^{\beta}$, where $n$ and $L$ represent the duration time of the protocol and the size of the probe, respectively, with $\alpha$ and $\beta$ as the corresponding exponents. While, without quantum effects, at best one can achieve $\alpha{=}\beta{=}1$, exploiting quantum features, such as entanglement and squeezing, enhances the scaling to $\alpha{>}1$ and $\beta{>}1$. A notable case is when $\alpha{=}2$ and $\beta{=}2$, which is called Heisenberg scaling. The most common way to capture this scaling is by initializing a dynamical protocol in maximally entangled states, such as Greenberger–Horne–Zeilinger (GHZ) states, and encoding the desired parameter through a unitary evolution, while the interaction between the system’s components remains negligible [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Alternatively, strongly correlated many-body systems can also be used a s a resource for achieving quantum enhanced sensitivity by harnessing a variety of quantum features such as criticality [118]. In this type quantum sensors, unlike GHZ-type sensors, the interaction between the components of the probe plays a crucial role in obtaining quantum enhanced sensitivity. Nonetheless, the obtainable precision is fundamentally bounded by the interaction details in the system [5, 6, 7].
@@ III Model
Figure 1: (a) Schematic for the DTC sensor introduced here. (b) Imbalance in the total magnetization of the chains $a$ and $b$ denoted as $\mathcal{I}(nT)$, versus the amplitude of the periodic-field $h_{a}$ at different stroboscopic times in a system of size L=6L{=}6. (c) Imbalance at even stroboscopic times in systems with various sizes, when $h_{a}{=}0.25$ and the system operates in the non-DTC region. In both panels, we set $\varepsilon{=}0.1$.
@@
shows persistent coherent oscillation between $\mathcal{I}(nT){=}1$ at even $n$’s obtained for $|\psi(0)\rangle$, and $\mathcal{I}(nT){=}{-}1$ at odd $n$’s obtained for the inversion partner of the initial state $|\tilde{\psi}(0)\rangle{=}\Pi_{j=1}^{L}\sigma^{ax}_{j}{\otimes}\sigma^{bx}_{j}|\psi(0)\rangle$. The origin of this period-doubling oscillations that last for $\tau{\propto}e^{\alpha L}$ with $\alpha{=}\ln(\varepsilon^{-1.033})$ is the emergence of $\pi$-Paired Floquet states as $|\Phi_{\pm}\rangle{\cong}(|\psi(0)\rangle{\pm}|\tilde{\psi}(0)\rangle){/}\sqrt{2}$ with eigenvalues differing by $\pi$ [112]. The sensing capability of general DTCs for estimating a uniform periodic-field has been investigated in Ref. [117]. The analytical calculations reveal the potential of capturing the quantum enhancement with respect to the system size as $\propto L^{2}$. One can expect to obtain the same results in sensing a uniform magnetic field using the presented system in Eq. 4. This motivates us to investigate the capability for sensing a non-uniform field. To this end, we expose the chain $a$ to a gradient periodic-field in the $z$-direction as
$H_{ac}(t)=B_{a}(t)G^{az}=h_{a}\sin(2\pi f_{ac}t)\sum_{j=1}^{L}j\sigma_{j}^{az},$ (6)
where $h_{a}$ and $f_{ac}$ are the amplitude and frequency of the periodic-field, respectively. See Fig. 1(a) for a system schematic. To trace the effect of this periodic-field on dynamics, we approximated $B_{a}(t)$ with a square-pulse form of the same period and alternating sign. This approximation is justified in our case because our focus is on the accumulated phase over time rather than the detailed structure of the continuous driving field. This approximation captures the essential contribution of the periodic-field to the dynamics. Under this assumption, $H_{\text{ac}}(t)$ becomes a piecewise static Hamiltonian that flips sign every period $T_{ac}$. For a periodic-field that resonates with the DTC, namely $T_{ac}{=}2T$, the dynamics can be well described by an effective Hamiltonian as
$H_{\mathrm{eff}}(t)=\begin{cases}H_{a}+H_{b}+\Theta_{n}G^{az}&(n{-}1)T<t\leqslant(n{-}\frac{1}{2})T\\ H_{I}+\Theta_{n}G^{az}&(n{-}\frac{1}{2})T<t\leqslant nT\end{cases}$ (7)
with $\Theta_{n}{=}h_{a}{\int_{(n-1)T}^{(n-\frac{1}{2})T}\sin(\frac{\pi\tau}{T})d\tau}{=}h_{a}\int_{(n-\frac{1}{2})T}^{nT}\sin(\frac{\pi\tau}{T})d\tau{=}(-1)^{n+1}\frac{h_{a}}{\pi J_{z}}$, as the time-averaged strength of the periodic-field. This allows us to express the Floquet operator as
$U_{n}{\simeq}e^{-i(H_{I}+\Theta_{n}G^{az})}e^{-i((H_{a}+H_{b})+\Theta_{n}G^{az})}.$ (8)
@@
Here we set $J_{z}{=}1$ as the unit of energy. During the first period, $t\in[0,T]$, the spins in the chain $a$ are primarily in $|{\uparrow,\cdots,\uparrow}\rangle_{a}$, and the field is positive. The accumulated phase during this interval is therefore positive. During the second period, $t\in[T,2T]$, the field becomes negative, but the spins flip to $|{\downarrow,\cdots,\downarrow}\rangle_{a}$, leading again to a positive phase accumulation. To show the effect of the implemented periodic-field on the DTC order, Fig. 1(b) plots the imbalance $\mathcal{I}(nT)$ versus the amplitude of the periodic-field $h_{a}$ over tens of stroboscopic times $n$ when the system initialized in $|\psi(0)\rangle$. By increasing $h_{a}$, the perfect and stable oscillations of the imbalance in stroboscopic times, namely $\mathcal{I}(2nT){=}1$, are replaced by nontrivial oscillations. This hints that our system goes through a sharp phase transition from a stable DTC order (for $h_{a}{<}h_{a}^{\max}$) to a region with no spontaneous breaking of DTTS (for $h_{a}{>}h_{a}^{\max}$), denoted as non-DTC region. Surprisingly, this transition, happening in a specific value denoted by $h_{a}{=}h_{a}^{\max}$ (dashed line), repeats in all the stroboscopic times, and, as we show later, is robust against the deviation parameter $\varepsilon$. In Fig. 1(c), we plot the dynamic of $\mathcal{I}(2nT)$ when $h_{a}{=}0.25$ and hence, the systems of different sizes operate in the non-DTC region. Clearly, by enlarging the system size, the amplitude of the nontrivial oscillations decreases considerably. This implies that, in systems of large enough sizes, these oscillations practically vanish even in the early time windows, signaling the thermalization of the system.
@@
In the following, we first show that the DTC phase can serve as a valuable resource for quantum-enhanced sensitivity and provide ultraprecise periodic-field sensing. Then, we complete our study by providing a robustness analysis of this enhancement to the perturbations in initialization, frequency offset, and crosstalk. We also explore the effect of $\varepsilon$ on the capability of our probe and propose an experimental protocol for its implementation.
@@ IV DTC sensor
Figure 2: (a) The QFI as a function of $h_{a}$ over different stroboscopic times. (b) The dynamical behavior of the QFI versus stroboscopic time $n$ at different values of $h_{a}$. In both panels (a) and (b), we set L=7L{=}7 and $\varepsilon{=}0.1$. (c) QFI as a function of $h_{a}$ for systems of different sizes over n=10n{=}10 period cycles when $\varepsilon{=}0.1$. (d) The values of QFI inside the DTC phase, namely $h_{a}{=}10^{-5}$, and in the transition point, namely $h_{a}{=}h_{a}^{\max}$, as a function of the system size $L$. The markers are the numerical results, and the solid lines provide the best fitting function as $\mathcal{F}_{Q}{\propto}L^{\beta}$ with $\beta{=}4.0902$ and $\beta{=}2.5887$ for a sensor in the DTC phase and in its transition point. The inset is the transition point $h_{a}^{\max}$ as a function of $L$. The best fitting function is obtained as $h_{a}^{\max}{\propto}L^{-1.0916}$.
To assess the performance of our DTC sensor for measuring the amplitude of the periodic-field, one can show that $U_{t}{=}\mathcal{T}e^{-i\int_{0}^{nT}(H(t)+H_{ac}(t))dt}{=}\Pi_{k=1}^{n}U_{k}$ for t=nTt{=}nT while $U_{k}$ is the propagator for one period cycle given in Eq. (8).
For a resonant periodic-field with the DTC, straightforward calculations lead to
$\partial_{h_{a}}U_{t}$ $={-i}U_{t}\sum_{k=1}^{n}\Big(\Pi_{l=1}^{k}U_{l}\Big)^{\dagger}\Big(\partial_{h_{a}}\Theta_{k}G^{az}\Big)\Big(\Pi_{l=1}^{k}U_{l}\Big)\leqslant iU_{t}nG^{az}/\pi.$ (9)
@@
[…]
wherein $D_{k}^{az}{=}\sum_{j=1}^{L}ja_{j}^{k}\sigma^{az}_{j}{\leqslant}\sum_{j=1}^{L}j\sigma^{az}_{j}{=}G^{az}$ and
$a_{j}^{k}{=}\frac{j^{2}\Theta_{k}^{2}+J_{ab}^{2}\cos^{2}\big(\sqrt{j^{2}\Theta_{k}^{2}+J_{ab}^{2}}\big)}{j^{2}\Theta_{k}^{2}+J_{ab}^{2}}.$ (12)
Using
@@
$4$ $\left(\langle\psi(0)|(\partial_{h_{a}}U_{t})^{\dagger}(\partial_{h_{a}}U_{t})|\psi(0)\rangle{-}|\langle\psi(0)|U_{t}^{\dagger}(\partial_{h_{a}}U_{t})|\psi(0)\rangle|^{2}\right)$
@@
and Eq. 9, one can simply obtain $\mathcal{F}_{Q}(t){\leqslant}4n^{2}(\Delta G^{az})^{2}/{\pi^{2}}$ in which $\Delta G^{az}{=}\sqrt{\langle(G^{az})^{2}\rangle{-}\langle G^{az}\rangle^{2}}$ stands for the variance. This bound depends on the variance of $G^{az}$ in the initial state. If the initial state has vanishing variance with respect to this generator, i.e. $\Delta G^{az}=0$, then the QFI bound becomes zero. This outcome is consistent with the fact that, in such a case, the state is an eigenstate of $G^{az}$. As a consequence, the dynamics generated by $G^{az}$ cannot imprint any distinguishable phase information on the state, and thus no sensitivity can be gained, leading to a vanishing QFI throughout the evolution. However, one can use the upper bound of the variance, defined by the square of the spectral norm defined as $G^{az}||^{2}{=}(\lambda_{\max}{-}\lambda_{\min})^{2}/4$ with $\lambda_{\max}{=}L(L{+}1)/2$ and $\lambda_{\min}{=}{-}L(L{+}1)/2$ as the largest and smallest eigenvalues of $G^{az}$. This results in $\mathcal{F}_{Q}{\leqslant}n^{2}L^{2}(L+1)^{2}/\pi^{2}.$ Capturing this upper bound that determines the ultimate precision in estimating $h_{a}$ requires satisfying two conditions. First, $a_{j}^{k}{\simeq}1$. Concerning Eq. 12, one see that $\varepsilon{=}0$ and $h_{a}{\ll}\sqrt{2}\pi^{2}/L$ can lead to $\cos^{2}\Big(\sqrt{j^{2}\Theta_{k}^{2}+J_{ab}^{2}}\Big){\simeq}1$. Second, the probe is to be initialized in the balanced superposition of the eigenstates corresponding to the extremal eigenvalues of $G^{az}$, namely $|\lambda_{\max}\rangle{=}|{\uparrow,\cdots,\uparrow}\rangle_{a}$ and $|\lambda_{\min}\rangle{=}|{\downarrow,\cdots,\downarrow}\rangle_{a}$. Therefor, the optimal initial state is $\frac{1}{\sqrt{2}}\big(|{\uparrow,\cdots,\uparrow}\rangle_{a}{\otimes}|{\downarrow,\cdots,\downarrow}\rangle_{b}{\pm}|{\downarrow,\cdots,\downarrow}\rangle_{a}{\otimes}|{\uparrow,\cdots,\uparrow}\rangle_{b}\big)$, concerning the conservation of the total magnetization in our protocol. In the following, we show that, although initialized in the product state $|\psi(0)\rangle$, achieving this ultimate precision even for $\varepsilon{\neq}0$ is permitted in our DTC probe, due to the freely emerged ”Schrödinger cat” state $(|\psi(0)\rangle{\pm}|\tilde{\psi}(0)\rangle){/}\sqrt{2}$ in the spectrum of DTC.